Let X~U(5,21) , calculate the expected value for this uniform distribution.


1
Expert's answer
2021-01-25T12:32:12-0500

The probability density function of X is f(x) = 1/(21-5) = 1/16.

The expected value of X is "E[X] = \int_5^{21}xf(x)dx = \int_5^{21}\frac{x}{16} dx =\frac{x^2}{32}|_5^{21} =(21^2-5^2)/32 = 13"


Answer. E[X]=13